# AutoProSAM: Automated Prompting SAM for 3D Multi-Organ Segmentation

### v3

@@ 6 Conclusion
In this paper, we introduced the AutoProSAM, a novel approach designed to enhance SAM for 3D multi-organ medical image segmentation. By extending SAM from 2D natural images to 3D medical images, our method addresses domain gaps and spatial disparities through parameter-efficient fine-tuning and an Auto Prompt Generator (APG) that automates prompt creation, eliminating the need for manual input. Extensive experiments on public and private CT-based datasets demonstrated the AutoProSAM’s superior performance over state-of-the-art models, evidenced by higher Dice and NSD scores. Our ablation studies also highlighted the importance of the APG and Multi-Layer Aggregation Mechanism (MLAM) in improving segmentation accuracy.
Limitations. While our AutoProSAM achieved superior performance in multi-organ segmentation tasks, certain limitations should also be noted. Our approach does not consider the pre-training of the image encoder via collecting a huge amount of medical images, nor does it focus on creating a universally adaptable model for multiple different datasets. Further related research would be worthwhile.

### v4

@@ 6 Conclusion
In this paper, we introduced AutoProSAM, a novel approach to enhance SAM for 3D multi-organ medical image segmentation. By adapting SAM from 2D natural images to 3D medical images, our method addresses domain gaps and spatial disparities through parameter-efficient fine-tuning and an Auto Prompt Generator (APG) that automates prompt creation, removing manual input. Extensive experiments on public and private CT-based datasets demonstrated AutoProSAM’s superior performance over state-of-the-art models, with higher Dice and NSD scores. Ablation studies also confirmed the critical role of APG and the Multi-Layer Aggregation Mechanism (MLAM) in boosting segmentation accuracy.